OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4 a. 11y + 4z = 6 2y + 5z = 3 b. y – 2z = 6 –8y – z = 3 c. –9y + z = –3 –y + 2z = –6 d. 3y + 3z = –3 11y + 4z = –6

OpenStudy (campbell_st):

option C the 3 equations x + 6y + z = 2 (1) x - 3y + 2z = -1 (2) x +5y +3z = -4 (3) Equation (2) - (1) Equation (3) - (1) x - 3y + 2z = -1 x + 5y + 3z = -4 - ( x + 6y + z = 2) - (x + 6y + z = 2) ----------------- ---------------- -9y + z = -3 -y + 2z = -6

OpenStudy (campbell_st):

hope it helps